You are driving a car with a speed of 60Km per hour. Your friend was driving a car with a speed of 80Km
per hour. Both of you have started at the same time how far your friend will be after 5 Hours and 30
minutes.

Answer by oberobic(2304)

d = r*t is the basic distance equation.
.
t = 5 hr 30 min
t = 5.5 hr
.
r = 60 km/hr
5.5 hr * 60 km/hr = 5.5*60 km
5.5*60 = 330 km
.
r = 80 km/hr
5.5 hr * 80 km/hr = 5.5*80 km
5.5*80 = 440 km
.
So your friend driving at 80 km/hr will be 110 km ahead.
.
You also could have just done one solution using the speed difference.
80 km/hr = 60 km/hr = 20 km/hr
.
5.5 hr * 20 km/hr = 110 km
.
Done.